### v1

# Linguistic Neuron Overlap Patterns to Facilitate Cross-lingual Transfer on Low-resource Languages

Abstract
The current Large Language Models (LLMs) face significant challenges in improving performance on low-resource languages and urgently need data-efficient methods without costly fine-tuning. From the perspective of language-bridge, we propose BridgeX-ICL, a simple yet effective method to improve zero-shot Cross-lingual In-Context Learning (X-ICL) for low-resource languages. Unlike existing works focusing on language-specific neurons, BridgeX-ICL explores whether sharing neurons can improve cross-lingual performance in LLMs or not. We construct neuron probe data from the ground-truth MUSE bilingual dictionaries, and define a subset of language overlap neurons accordingly, to ensure full activation of these anchored neurons. Subsequently, we propose an HSIC-based metric to quantify LLMs-internal linguistic spectrum based on overlap neurons, which guides optimal bridge selection. The experiments conducted on $2$ cross-lingual tasks and $15$ language pairs from $7$ diverse families (covering both high-low and moderate-low pairs) validate the effectiveness of BridgeX-ICL and offer empirical insights into the underlying multilingual mechanisms of LLMs.
1 Introduction
Although Large Language Models (LLMs) have demonstrated impressive multilingual capacities, there is still significant space for improving the performance on low-resource languages Huang et al. (2024); Nazi et al. (2025). To address this issue, especially avoiding costly post-training Muller et al. (2021); Yong et al. (2023), it is critical to fully investigate the multilingual understanding and transferring ability in LLMs.
Recent research has increasingly focused on data-efficient methods, particularly Cross-lingual In-Context Learning (X-ICL) Winata et al. (2021); Tanwar et al. (2023); Nazi et al. (2025); Cahyawijaya et al. (2024), which surprisingly works well on low-resource languages, primarily due to LLMs are in-context low-resource language learners Brown et al. (2020b); Cahyawijaya et al. (2024). For instance, in the Arabic-to-Hebrew Bilingual Lexicon Induction (BLI) task, the zero-shot baseline accuracy in LLaMA 3 is 47.0%. However, simply specifying English as a bridge language in a zero-shot setting boosts accuracy to 64.5%, which significantly outperforms even the two-shot X-ICL. This observation motivates us to further explore: How can we improve cross-lingual capabilities of LLMs on low-resource languages by selecting an optimal bridge language in X-ICL? Should the selection be purely data-driven, favoring high-resource bridge languages Vulic et al. (2020)? Or can human linguistic knowledge, such as language genealogy, or established evolutionary taxonomies, offer a more effective alternative Stanczak et al. (2022); Wang et al. (2024)?
To fully investigate this issue from a systematic perspective, we leverage linguistic neuron Tang et al. (2024) to guide optimal bridge language selection in X-ICL. However, there are two limitations when applying neuron-based interpretation Cao et al. (2024); Tang et al. (2024); Liu et al. (2024) on low-resource languages:
$\bullet$ Inaccurate Neuron Activation. Current work often relies on multilingual corpora like Wikipedia Foundation (2024) to probe internal-neurons, without verifying whether LLMs truly understand the input. This may lead to unreliable neuron activations, particularly for low-resource languages. When LLMs poorly understand the probe input, they may instead activate neurons largely for processing unfamiliar or noisy input.
$\bullet$ Lacking guidance for cross-lingual transfer. Current work focuses on analyzing the distribution of neurons Stanczak et al. (2022); Tang et al. (2024); Liu et al. (2024), yet how to leverage internal neurons to enhance cross-lingual transfer remains underexplored. Recent work argues that language-specific neurons do not facilitate cross-lingual transfer Mondal et al. (2025). This raises a critical question: Whether sharing neurons can improve cross-lingual transfer in LLMs? This exploration is also important to transfer language neuron research to actionable strategies for enhancing LLMs’ multilinguality.
Motivated by this, we propose a simple yet effective bridge method, BridgeX-ICL, to improve LLMs’ cross-lingual capabilities, especially on low-resource languages. To address the inaccurate activation issue, we construct probe data by leveraging the ground-truth bilingual lexicon MUSE Conneau et al. (2017). We then collect bilingual word pairs that LLMs can translate accurately and prompt them to LLMs using answer generation in bi-directions. To address the cross-lingual guidance issue, we first explore overlap neurons’ features and their impact on cross-lingual transfer, and then propose a bridge selection strategy based on HSIC Gretton et al. (2005). Furthermore, we measure the linguistic spectrum in LLMs based on overlap neurons and compare it with human language genealogy from Glottolog Trees Hammarström et al. (2023). We conduct extensive experiments on $2$ cross-lingual tasks and $15$ language pairs from $7$ diverse families. Our main contributions and findings are as follows:
  • To the best of our knowledge, this is the first work to explore language-bridge for zero-shot X-ICL to improve LLMs’ performance on low-resource languages.
  • We construct neuron probe data and use them to fully activate the anchored overlap neurons. We also propose a HSIC-based metric to quantify the similarity between overlapping-neurons and specific neurons for making optimal bridge selection in X-ICL.
  • We validate the efficacy and generalization of BridgeX-ICL on 2 cross-lingual tasks and $15$ language pairs. Here are empirical findings: 1) Strong neural overlaps align with human linguistic taxonomy within language families, but do not consistently hold across families. 2) Overlap neurons embody shared semantic information no matter among languages within or across families. 3) BridgeX-ICL improves the performance on 2 cross-lingual tasks across 15 language pairs by an average of 6.02% and 5.25% over zero-shot baselines.
Figure 1: An illustration of BridgeX-ICL approach, consisting of three steps: Neuron probe data construction; Language neurons and their overlappings detection; Optimal bridge $L_{*}$ selection based on HSIC similarity.
2 Related Work
2.1 Cross-lingual In-context Learning
LLMs face significant challenges when applied to low-resource languages Costa-jussà et al. (2022); Muennighoff et al. (2023); Huang et al. (2024), mainly due to insufficient training data and the curse of multilinguality Conneau et al. (2020). To address these issues without updating model parameters, Cross-lingual In-context Learning (X-ICL), an extension of in-context learning (ICL), has recently gained attention Brown et al. (2020a). Prior studies Winata et al. (2021); Tanwar et al. (2023); Nazi et al. (2025); Cahyawijaya et al. (2024) have demonstrated that LLMs act as effective few-shot multilingual learners, with few-shot ICL even outperforming fine-tuned language-specific models on several tasks Winata et al. (2021). However, few-shot X-ICL’s performance is highly dependent on the context and the selection of examples, especially for unconventional or ambiguous languages Philippy et al. (2023); Nazi et al. (2025). Consequently, existing research mainly focuses on optimizing few-shot example selection. To the best of our knowledge, we are the first to explore X-ICL explicitly from the perspective of leveraging language bridges.
2.2 Linguistic Neuron in LLMs
Recent research Stanczak et al. (2022); Tang et al. (2024); Liu et al. (2024); Cao et al. (2024); Wang et al. (2024) has revealed that language-related neurons exist in FFN layers of transformer architecture. Deactivating these neurons will play a vital impact on LLMs’ multilingual capacities. Beyond uncovering multilingual mechanisms, some research has gone to explore the neuron pattern across languages Wang et al. (2024); Stanczak et al. (2022) and its impact on cross-lingual performance Mondal et al. (2025); Zhang et al. (2025). Specifically, Wang et al. (2024) observed that similar languages may not exhibit significant neuron sharing in LLMs like BLOOM, suggesting that neuron sharing does not fully align with language similarity. Furthermore, recent work argues that language-specific neurons do not facilitate cross-lingual transfer Mondal et al. (2025). This raises a critical question: Whether sharing neurons can improve cross-lingual performance in LLMs. Motivated by these findings, we aim to further investigate LLM-internal neuron sharing across languages and its impact. In particular, we define a subset of language-overlapping neurons and explore whether they can serve as internal bridges to support cross-lingual inference.
3 Methodology
3.1 Task Statement
Given language set $\mathcal{L}=\{L_{1},...L_{|\mathcal{L}|}\}$, this work aims to measure the linguistic genealogy learned by LLMs from language-overlapping neurons and then use the quantified linguistic similarity to guide the bridge language selection in X-ICL.
Figure 1 depicts three main steps of our approach: ① Neuron probe data construction; ② Language neurons and their overlappings detection; ③ Bridge language selection based on a modified HSIC dependency estimation to measure linguistic distance, selecting $L_{*}$ from the candidate set $L_{\text{candidate}}$ to facilitate X-ICL from source language $L_{s}$ to target language $L_{t}$.
3.2 Probe Data Construction
We utilize two types of probe data for language neuron identification and optimal bridge selection. The former focuses on language-specific neurons and can use existing multilingual corpora. In our work, we adopt FLORES+ NLLB Team et al. (2024), a high-quality parallel corpus released by Meta, and combine its devtest and test sets to obtain 2,000 parallel sentences for each language.
Bridge selection for X-ICL needs to consider both language-specific neurons and those involved in cross-lingual tasks. Inspired by findings in task-specific neurons Song et al. (2024), we hypothesize certain neurons are associated with cross-lingual transfer whose manipulation and measurement should not rely on feeding monolingual input. Therefore, we construct probe data by leveraging bilingual word translations.
Prompt Design We collect $d$ (i.e., $100$) word pairs that LLMs can translate accurately per language pair. We prompt these word pairs to LLMs in both directions of $L_{1}$-$L_{2}$ and $L_{2}$-$L_{1}$, ensuring neurons linked to $L_{1}$ and $L_{2}$ are fully activated. Instead of feeding word pairs directly, we prompt LLMs to generate translations, which guarantees accurate and pronounced neuron activation. Examples of probe data for $3$ language pairs are shown below.
3.3 Linguistic Overlap Neurons
3.3.1 Neurons in LLMs
Neurons identification is based on Tang et al. (2024), which hypothesizes that language neurons are mainly located in the Feed-Forward Network (FFN) layers. Given the transformation in the $i$-th layer:
$\boldsymbol{h}_{i}=\sigma(\tilde{\boldsymbol{h}_{i}}\boldsymbol{W}_{1}^{i})\cdot\boldsymbol{W}_{2}^{i}$ (1)
where $\tilde{\boldsymbol{h}_{i}}$ is the hidden state input to the $i$-th layer and $\sigma(\cdot)$ denotes the activation function. $\boldsymbol{W}_{1}^{i}\in\mathbb{R}^{d\times N}$ and $\boldsymbol{W}_{2}^{i}\in\mathbb{R}^{N\times d}$ are the learned parameters. Here, a neuron is defined as a linear transformation of a single column in $\boldsymbol{W}_{1}^{i}$ and there are $N$ neurons in each layer. The activation value of the $j$-th neuron is $\sigma(\tilde{\boldsymbol{h}_{i}}\boldsymbol{W}_{1}^{i})_{j}$. If this value exceeds 0, the neuron is considered activated.
3.3.2 Overlap Neuron Identification
First, we identify neurons $\mathcal{T}_{k}$ associated with specific languages $L_{k}$. Unlike existing work Wang et al. (2024); Mondal et al. (2025) using LAPE Tang et al. (2024) to identify neurons with high activation probability for one language but low for others, which is less effective to detect neuron relationships across languages, we identify neurons $\mathcal{T}_{k}$ for each language $L_{k}$ based their activation frequency. Let $f_{k,j}$ denote the activation frequency of neuron $n_{j}$ when processing tokens of language $L_{k}$. Neurons with the top $\tau\cdot N$ activation frequencies are selected into $\mathcal{T}_{k}$ based on a threshold $\tau$.
Overlap Neuron Definition. For languages $L_{u}$, $L_{v}$, given their associated neurons $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$, the overlap neurons are the interaction of $\mathcal{T}_{u}$ and $\mathcal{T}_{v}$. For the $i$-th layer, we have $\mathcal{T}_{u,v}(i)=\mathcal{T}_{u}(i)\cap\mathcal{T}_{v}(i)$.
Linguistic Neuron Similarity. We measure the linguistic similarity between $L_{u}$ and $L_{v}$ based on their overlap neurons’ activation frequency as:
$\text{sim}(\mathcal{T}_{u},\mathcal{T}_{v})=\frac{\boldsymbol{f_{u}}\cdot\boldsymbol{f_{v}}}{\lVert\boldsymbol{f_{u}}\rVert\lVert\boldsymbol{f_{v}}\rVert}$ (2)
where $f_{u,j}\in\boldsymbol{f_{u}}$ and $f_{v,j}\in\boldsymbol{f_{v}}$ denote the activation frequency of the $j$-th neuron in $\mathcal{T}_{u,v}$ when processing tokens from $L_{u}$ and $L_{v}$, respectively.
3.3.3 Overlap Neurons Patterns
Second, we use the constructed probe data to explore overlap neurons’ features and their generalized impact on cross-lingual transfer so that we can utilize them to guide bridge language selection. We make two observations:
Figure 2: Language-overlapping neurons on distant pair (Arabic-Swahili) and close pair (Arabic-Hebrew).
Figure 3: Layer-wise latent embeddings projected with MDS in French-Swahili translation. A rainbow-colored path traces the latent embeddings across 32 layers. The predicted Swahili tokens are in green and their correct English tokens in orange.
$\bullet$ Similar languages share more neurons than distant ones. For example, Arabic-Hebrew within the same language family has more overlapping-neurons than Arabic-Swahili across families, presented in Figure 2, This suggests the potential of neural overlap to measure language distance.
$\bullet$ Overlap neurons concentrate in middle-layers and final-layers and serve distinct roles of semantic understanding and language coding for next-token prediction. This is also evidenced by neurons deactivation in Figure 7. Neurons in final layers are task-related and are responsible for cross-lingual generation. To examine whether middle-neurons handle semantic understanding, we employ a technique called logit lens Nostalgebraist (2020) to visualize the latent embeddings passing through neuron layers. We test in French-Swahili translation and select 60 word pairs that LLaMA 3 can translate accurately. After recording the model’s latent embeddings at each layer for next-token prediction, we use classical multidimensional scaling (MDS) to embed them in a 2D space, presented in Figure 3. The embedding trajectory is marked in a rainbow-colored path (e.g., red = layers 1-4, violet = layers 25-32). We can observe French inputs and the correct English next tokens cluster in middle layers, suggesting LLMs rely on the knowledge in high-resource languages like English to perform cross-lingual reasoning. Neurons in middle-layers should be prioritized over final-layers when measuring language similarity.
3.4 Bridge Language Selection
Based on the above observations, we quantify activation similarity between source-target overlap neurons and bridge-specific neurons and identify the optimal bridge language to facilitate X-ICL.
Given a language pair $L_{s}$ and $L_{t}$ and their overlap neurons $\mathcal{T}_{s,t}$, we obtain the activation value matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{T}_{s,t}|\times 2d}$ when probing LLMs with $d$ samples in both directions for balanced activation in $L_{s}$ and $L_{t}$. Simultaneously, we obtain $\mathbf{Y}\in\mathbb{R}^{|\bar{\mathcal{T}_{y}|}\times 2d}$, the activation value matrix for bridge language $L_{y}\in\mathcal{L}_{\text{candidate}}$, and $\bar{\mathcal{T}_{y}}=\mathcal{T}_{y}-\mathcal{T}_{s,t}-\mathcal{T}_{y^{\prime}}$ represents language-specific neurons in $L_{y}$ excludes neurons in $\mathcal{T}_{s,t}$ and $\mathcal{T}_{y^{\prime}}$, where $L_{y^{\prime}}\neq L_{y}$.
We employ the Hilbert-Schmidt Independence Criterion (HSIC) Gretton et al. (2005) to measure the nonlinear dependency between activation matrices of $\mathbf{X}$ and $\mathbf{Y}$. Average pooling will be performed to standardize matrices of $\mathbf{X}$ and $\mathbf{Y}$ to have the same row dimension $n$. The formal HSIC is calculated as: $\text{HSIC}(\mathbf{X},\mathbf{Y})=n^{-2}\text{Tr}(\mathbf{KHLH})$, where $\mathrm{Tr}(\cdot)$ is the trace operation, $\mathbf{K},\mathbf{L}\in\mathbb{R}^{n\times n}$ are learned kernel matrices for $\mathbf{X}$ and $\mathbf{Y}$. $\mathbf{H}=\mathbf{I}_{n\times n}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$ is a centering matrix, where $\mathbf{I}_{n\times n}$ is the identity matrix of size $n\times n$, $\mathbf{1}_{n}$ is a vector of $n$ ones. Rather than computing HSIC over the entire activation matrices, we adopt a bidirectional maximum matching strategy to measure the strongest dependency between individual neurons in one set and the entire distribution of the other, computed as:
$\mathrm{HSIC}(\bar{\mathcal{T}}_{y},\mathcal{T}_{s,t})=$ (3)
$\frac{1}{2}\left(\max_{i}\mathrm{HSIC}(\mathbf{x}_{i},\mathbf{Y})+\max_{j}\mathrm{HSIC}(\mathbf{X},\mathbf{y}_{j})\right)$
where $\mathbf{x}_{i}\in\mathbf{X},\mathbf{y}_{j}\in\mathbf{Y}$. We compute the dependency scores per layer, then average them across middle $K$ layers to get $L_{y}$’s selection probability:
$p(L_{y}|L_{s}\text{-}L_{t})=\frac{1}{K}\sum_{i=1}^{K}\text{HSIC}\left(\bar{\mathcal{T}}_{y}(i),\mathcal{T}_{s,t}(i)\right)$ (4)
where $K$ is determined according to embedding semantic similarity and discussed in section B.3. Finally, the optimal bridge $L^{*}$ is selected by:
$L^{*}=\arg\max_{L_{y}\in\mathcal{L}_{\text{candidate}}}p(L_{y}|L_{s}\text{-}L_{t})$ (5)
4 Experiment
4.1 Experiment Setup
Implementation. We evaluate BridgeX-ICL on $2$ cross-lingual tasks and $15$ languages covering $7$ diverse language families: Indo-European: English (En), German (De), French (Fr), Italian (It), Portuguese (Pt), Spanish (Es); Uralic: Finnish (Fi), Hungarian (Hu); Afro-Asiatic: Arabic (Ar), Hebrew (He); Austronesian: Indonesian (Id), Tagalog (Tl); Sino-Tibetan: Chinese (Zh); Japonic: Japanese (Ja); Niger-Congo: Swahili (Sw).
The evaluation focuses on LLMs’ cross-lingual transfer on low-resource languages. Therefore, we take He, Tl, Sw, and Ja as target languages to build $15$ cross-lingual pairs, covering moderate-to-low (e.g., Ar-Sw), and high-to-low (e.g., En-He) pairs, both within and across families. The classification of high-, moderate-, and low-resource languages is based on their proportion in LLM’s training corpora, following previous work Cieri et al. (2016). Since the bridge language should be well supported by LLMs, we take 6 languages in Indo-European family as candidate bridges.
Datasets. To evaluate the generalization of bridge selection beyond the BLI task, we also evaluate on an additional cross-lingual task: Machine Reading Comprehension (MRC) using the Belebele dataset Bandarkar et al. (2024).
To evaluate low-resource languages, a key challenge lies in lacking evaluation benchmarks. Although the ground truth MUSE Conneau et al. (2017) provides $110$ bilingual dictionaries for BLI task, it does not cover the tested 15 language pairs. To solve this, we used English as a pivot to build $L_{s}$-$L_{t}$ dictionary from $L_{s}$-English and $L_{t}$-English. For languages not in MUSE (e.g., Swahili), we extracted word pairs from wiktionary_bli Izbicki (2022) to build En-Sw. We checked all word pairs using both Google and Microsoft translators to ensure quality and selected $1,000$ word pairs for each language pair that are consistently validated by both systems. The constructed BLI dictionaries are available at: https://anonymous.4open.science/r/BLI--0481/.
Metrics. For BLI task, we use Precision@N metric, which measures the accuracy of the model’s top $N$ candidate translations. In this study, $N$ is set to 1. For the MRC task, we use accuracy to measure whether the model selects the correct answer from multiple choices.
LLMs. We conducted experiments on two open-source LLMs: LLaMA-3-8B Grattafiori et al. (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023). Their training corpora cover 176 and 53 languages, respectively, which include all the experimental low-resource languages and allow us to explore the underlying linguistic mechanisms.
[…]
4.2 Main Results
4.2.1 LLMs’ Linguistic Spectrum Discussion
(a) LLaMA 3’s Linguistic spectrum
(b) Human language similarity from Glottolog
Figure 4: Comparison of linguistic spectrum calculated based on overlap neurons in LLaMA 3 and human language similarity derived from Glottolog Phylogenetic Trees, including 15 languages from 7 families. Darker blue indicates a higher language similarity.
This section discusses the linguistic similarities across $15$ languages from $7$ families, calculated based on overlap neurons in LLaMA 3 and Mistral, as presented in Figure 4 (a) and Figure 8 in B.2, respectively. To evaluate how closely the linguistic spectrum learned by LLMs align with that of human languages, we leverage Glottolog Phylogenetic Trees Hammarström et al. (2023), which encode hierarchical relationships among 8,000+ human languages, to draw human language similarity in Figure 4(b). The detailed linguistic similarity computed based on Glottolog is in Appendix A. In Figure 4, color intensity represents degree of similarity between languages, with darker blue indicating stronger degree and the diagonal is self-similarity (1.0).
Table 1: Comparison of BLI task improvement on 15 language pairs. The highest gains are marked with bold in few-shot and zero-shot with bridge methods. ’-’ indicates the selected bridge is either the source or target language.
LLaMA-3-8B
Method Zh-Ja Zh-He Zh-Tl Zh-Sw Ar-Ja Ar-He Ar-Tl Ar-Sw Id-Ja Id-He Id-Tl Id-Sw En-He En-Tl En-Sw
Zero-shot 67.10 44.10 42.60 31.20 69.90 47.00 46.70 39.10 62.50 44.70 49.30 25.90 56.90 60.00 28.80
One-shot +3.20 +12.60 +1.20 +3.00 +4.20 +13.50 -4.80 +0.60 +7.40 +7.50 +0.70 +6.00 +18.60 -6.30 +4.40
Two-shot +9.40 +16.90 +2.20 +5.10 +9.40 +13.90 -0.30 +3.50 +16.00 +15.90 +5.90 +6.10 +23.90 -5.50 +6.50
Three-shot +6.70 +22.20 +3.70 +6.80 +7.80 +16.90 +1.50 +3.70 +16.70 +22.90 +10.10 -4.30 +26.30 -4.00 +6.90
Few-shot Four-shot +12.50 +20.50 +3.20 +7.00 +7.70 +15.80 +0.80 +3.20 +14.30 +22.00 +10.60 -6.00 +26.10 -3.70 +7.40
Human source +2.80 +6.70 +3.80 +1.50 -9.50 - -7.40 -0.60 -11.90 -3.20 - -3.90 +12.30 -3.30 -1.70
Human target +2.80 +9.60 -0.10 +5.00 -0.50 - -7.10 - -12.10 +3.80 - +1.20 +16.40 -2.30 +2.30
English bridge +10.80 +12.60 +11.40 +4.80 +10.70 +17.50 +10.10 +3.30 +3.60 +9.30 +10.40 +2.50 - - -
Sharing matter +9.50 +14.50 +8.40 +2.60 +6.40 +17.10 +6.10 +3.90 +2.60 +15.20 +7.70 +2.40 +12.30 -3.30 -1.70
IoU Score +10.80 +12.60 +11.40 +4.80 +10.70 +13.20 +5.30 +3.50 +2.30 +15.20 +7.40 +3.80 +9.70 -6.10 -1.80
LAPE_overlap +10.50 +13.90 +11.40 +3.00 +10.70 +17.50 +10.10 +3.30 +3.60 +9.30 +10.40 +2.50 +12.30 -6.10 -1.80
Zero-shot
with bridge
Ours +10.80 +12.60 +11.40 +4.80 +10.70 +17.50 +10.10 +3.30 +3.60 +16.60 +11.70 +4.10 +14.90 -1.30 -2.60
Mistral-7B
Method Zh-Ja Zh-He Zh-Tl Zh-Sw Ar-Ja Ar-He Ar-Tl Ar-Sw Id-Ja Id-He Id-Tl Id-Sw En-He En-Tl En-Sw
Zero-shot 57.80 26.20 34.10 8.40 52.50 32.30 28.00 9.10 48.40 36.20 40.60 8.60 47.80 45.70 8.20
One-shot -7.20 -0.40 -1.90 +1.60 -8.10 -7.40 -2.00 +2.10 +5.60 +2.50 -3.70 +1.20 +0.40 -13.80 +2.40
Two-shot +0.40 +0.80 -2.90 +2.10 -5.10 -7.00 -0.40 +1.60 +10.10 +3.20 -0.60 +1.60 +0.50 -3.90 +3.20
Three-shot +3.00 +0.70 -1.60 +2.20 -2.90 -6.50 +0.20 +2.00 +10.70 +3.30 +0.80 +2.00 0.00 -2.80 +3.00
Few-shot Four-shot +3.20 +1.20 -0.50 +2.60 -2.90 -6.60 0.00 +2.20 +10.70 +4.10 +2.50 +2.60 +1.20 -2.00 +2.50
Human source -3.60 -0.20 +0.90 +0.90 -8.70 - -2.20 +0.80 -5.70 -5.70 - -0.40 -0.10 +2.30 +2.10
Human target -3.60 -1.60 -0.60 +0.20 -9.50 - +4.90 - -4.70 -7.20 - -0.20 -9.00 +1.40 +1.20
English bridge +7.90 +8.60 +6.40 +1.20 +8.90 +2.70 +8.60 +1.20 +9.60 +2.20 +2.50 +0.20 - - -
Sharing matter +7.90 +8.60 +6.40 +1.20 +4.60 +2.70 +8.60 +1.20 +9.60 +2.20 +2.50 +0.20 -0.10 +1.50 +0.90
IoU Score +0.50 +5.30 +4.90 +2.20 +5.90 +1.10 +7.50 +1.60 +2.40 -2.80 -0.50 +1.00 -1.00 +1.50 +2.20
LAPE_overlap +0.50 +3.50 +2.20 +2.20 +8.90 +2.70 +7.60 +1.20 +9.60 +2.20 +1.00 +0.20 -0.10 +4.50 +2.00
Zero-shot
with bridge
Ours +7.90 +8.60 +6.40 +1.20 +8.90 +2.70 +8.60 +1.20 +2.40 -3.30 +3.30 +1.00 -0.10 +1.50 +2.00
Linguistic spectrum learned by LLMs are not fully aligned with human languages. We can observe a strong neural similarity within language families, marked using red text box in Figure 4, which matches human linguistic taxonomy. For example, within Afro-Asiatic family, both from Arabic (Ar) and Hebrew (He), have a high neuron overlap (0.470), greater than Arabic-Swahili with 0.161. In addition, high-resource Indo-European languages, such as Fr-It and Pt-Es, show the highest overlap scores, with darkest blue in the bottom-right corner of the heatmap. But this alignment breaks down between high- and low-resource languages, like Arabic-French with similarity $0.479$.
LLMs build their own distinct understanding of language relationships. The calculated linguistic spectrum of LLaMA 3 and Mistral are similar but not the same. The two models may choose different bridges for the same language pair, as discussed later. This is likely because their internal linguistic relationships are primarily shaped by the distribution of languages in training corpora, as noted in Philippy et al. (2023). This explains why Arabic has the strongest similarity ($0.479$) with French in Romance family, rather than with Hebrew ($0.470$) in the same Afro-Asiatic.
4.2.2 Cross-lingual Results Analysis
This section compares performance of BridgeX-ICL against various baselines on BLI task across $15$ language pairs, as presented in Table 1.
Table 2: Evaluation on MRC cross-lingual task. Red color highlights the different bridge selection and bold marks the highest gains at each language pair.
LLaMA-3-8B Mistral-7B
Bridge Zero-shot Ours Bridge Zero-shot Ours
Zh-Ja En 61.80 -0.40 En 66.20 +6.20
Zh-He En 56.00 +7.20 En 50.20 +10.20
Zh-Tl En 57.20 +3.00 En 60.60 +7.80
Zh-Sw En 48.60 +10.40 Pt 43.00 +1.00
Ar-Ja En 52.20 +3.20 En 48.40 +7.40
Ar-He En 51.20 +4.00 En 42.00 +5.20
Ar-Tl En 46.40 +7.20 En 47.60 +4.60
Ar-Sw En 40.60 +12.60 En 30.40 +9.80
Id-Ja En 56.20 +8.20 Pt 63.40 +2.00
Id-He Es 58.20 +7.00 Es 45.80 +6.40
Id-Tl Fr 58.40 +4.20 De 55.60 +3.00
Id-Sw Fr 49.80 +5.20 Pt 39.60 -2.80
En-He Es 70.60 +8.00 Es 61.20 +3.80
En-Tl Fr 72.80 +3.20 Es 72.20 +2.40
En-Sw Fr 64.60 +7.40 Fr 51.20 +0.20
We can find: 1) LLMs exhibit poor and imbalanced performance on low-resource languages. For example, LLaMA 3 achieves its best BLI performance of $69.90$ on Ar-Ja pair, but worst of $25.90$ on Id-Sw pair. 2) LLMs are few-shot multilingual learners. However, few-shot X-ICL does not consistently yield stable gains, e.g., one-shot falling below zero-shot, and will not lead to further gains when the number of shots goes beyond $3$. It indicates when applying few-shot X-ICL, 3-shots will be enough. 3) Zero-shot with bridge is a simple yet data-efficient strategy for low-resource languages. BridgeX-ICL finds $9$ optimal bridges out of $15$ language pairs, achieving average performance of two-shot X-ICL across all pairs, followed by the English-bridge method. While human source/target methods are the least effective. It seems using English as the default bridge is cost-effective, which will be discussed in section 5.1.
5 Discussion
5.1 Application of Bridge Language
Beyond BLI task, Table 2 evaluates BridgeX-ICL on the MRC cross-lingual task. The prompt for zero-shot with bridge in MRC is detailed in Appendix E. Results show our approach still works well on MRC task and benefits more on LLaMA 3 than Mistral. For example, BridgeX-ICL improves performance of LLaMA 3 by an average of 6.036.03% over the zero-shot baseline across 15 language pairs, while the average improvement is 4.484.48% on Mistral.
English is selected as the optimal bridge in 9 out of 15 language pairs. The reason is partly due to LLMs’ unbalanced language abilities across 5 candidate bridges. As discussed in Figure 3 another key factor is LLMs’s inherent preference of English-pivot during cross-lingual transfer.
5.2 Ablation Study
In this part, we conduct ablation study to evaluate the impact of neuron probe data and the proposed HSIC similarity metric on bridge selection, using the BLI task for example.
Table 3: The impact of neuron probe data. ‘w/o *’ denotes replacing our constructed probe data with bilingual tokens extracted from FLORES+ dataset.
LLaMA-3-8B Mistral-7B
w/o * Ours w/o * Ours
Zh-Ja 76.40 77.90 $\uparrow$ 63.40 65.70 $\uparrow$
Zh-He 56.70 56.70  32.70 34.80 $\uparrow$
Zh-Tl 51.10 54.00 $\uparrow$ 38.20 40.50 $\uparrow$
Zh-Sw 36.40 36.00 $\downarrow$ 10.60 9.60 $\downarrow$
Ar-Ja 77.80 80.60 $\uparrow$ 58.40 61.40 $\uparrow$
Ar-He 64.10 64.50 $\uparrow$ 33.40 35.00 $\uparrow$
Ar-Tl 52.00 56.80 $\uparrow$ 33.30 35.50 $\uparrow$
Ar-Sw 41.60 42.40 $\uparrow$ 10.70 10.70
Id-Ja 65.10 66.10 $\uparrow$ 50.80 58.00 $\uparrow$
Id-He 59.90 61.30 $\uparrow$ 33.40 38.40 $\uparrow$
Id-Tl 57.00 61.00 $\uparrow$ 40.20 40.20
Id-Sw 28.30 30.00 $\uparrow$ 9.20 9.20
En-He 66.60 71.80 $\uparrow$ 47.70 47.70
En-Tl 58.70 58.70  47.70 48.00 $\uparrow$
En-Sw 26.20 26.20  10.20 10.30 $\uparrow$
Table 3 presents the results of ablation experiments to compare “w/o *” with our constructed probe data, where “w/o *” denotes replacing the constructed probe data with the bilingual tokens extracted from FLORES+ NLLB Team et al. (2024). Appendix C illustrates the detailed construction for “w/o *”. These results highlight impact of probing data in neuron manipulation. Table 4 in Appdedix C compares the performance of HSIC and Cosine similarity and shows the proposed HSIC helps to capture the dependency between language-overlapping neurons and specific neurons.
5.3 Overlapping Neuron Distribution
Figure 5: Distribution of overlap neurons in language pairs within and across families in LLaMA 3.
Figure 6: Layer-wise activation frequency of neurons viewed from language families in LLaMA 3.
This section analyzes the distribution of overlap neurons. Figure 5 compares neurons in language pairs within the same family (e.g., Ar-He) and across families (e.g., Ar-Sw). Obviously, Ar-He shares more overlap neurons. Similar observations can be found in comparing language pairs from different source languages (Zh, Ar, Id, and En ) to a same target language He (Figure 10 in B.4) From the perspective of language families, Figure 6 examines the activated behaviors of neurons in low-resource languages. Obviously, low-resource languages within the Uralic family have the highest activation frequency, while Indo-European languages have the lowest. We hypothesize LLMs activate neurons more frequently for processing low-resource languages due to their perceived difficulty.
6 Conclusion
In this work, we explore whether sharing neurons can improve LLMs’ cross-lingual performance on low-resource languages. We propose a simple yet effective language-bridge approach with the help of neuron interpretation. To ensure accurately and fully activate overlap neurons across languages, we construct neurons probe data from the ground-truth MUSE dictionaries. By quantifying neuron similarity, we seek the optimal bridge for X-ICL and conduct extensive experiments to validate the efficacy and generalization of our approach.
Limitations
This work focuses on sharing neurons across languages and relies on evaluated data to validate the effectiveness of our approach. Due to the limitation of lacking evaluation benchmarks for low-resource languages, the experiments are conducted on 15 language pairs and have not been extensively validated on a wide range of low-resource languages. We select 4 typologically diverse low-resource languages from distinct families and test their performance on 2 cross-lingual tasks. Second, our study reveals that high-quality probe data is essential for accurately analyzing neurons’ behaviors of low-resource languages. The proposed linguistic distance measurement is probe-data-induced, providing qualitative insights, but lacks quantitative precision.
Acknowledgments
This work was supported by the National Social Science Foundation (No.24CYY107) and the Fundamental Research Funds for the Central Universities (No.2024TD001).
Appendix A Appendix: Linguistic Similarity Based on Glottolog Phylogenetic Trees
We leverage Glottolog verion 5.1 Hammarström et al. (2023) as a foundational phylogenetic framework to calculate the linguistic similarity of human languages. It has two key steps: data preprocessing and similarity calculation.
Data Preprocessing. The preprocessing pipeline consists of three steps: 1) Glottocode Identifiers with regex pattern matching to ensure unambiguous language node identification. 2) Standardize node naming with underscores (e.g., [sini1245]→ _sini1245_), ensuring consistent formatting in downstream phylogenetic analyses. 3) Mitigate encoding conflicts through temporary file caching. These steps preserve both accurate phylogenetic tree parsing and computational compatibility.
Similarity Calculation. The proposed metric integrates two well-established phylogenetic principles from historical linguistics, which includes node distance normalization and depth-adjusted compensation.
First, building upon Wichmann & Holman’s framework for typological stability assessment Wichmann and Holman (2009), we compute the inter-language distance $d(L_{1},L_{2})$ between languages $L_{1}$ and $L_{2}$ using ETE3’s optimized tree traversal algorithms Huerta-Cepas et al. (2016) and then normalize it to make distances comparable across language families, calculated as:
[…]
where $\hat{D}$ is the family-specific maximum. For example, $\hat{D}=80$ for Sino-Tibetan languages, reflecting their deep internal divergence, whereas $\hat{D}=75$ for Indo-European languages, due to their relatively shallower subgroup structure.
Second, depth-adjusted compensation aims to mitigate biases introduced by uneven tree depth and family-specific structural variation. Following the work Gray et al. (2009) to calculate depth disparity factor $\delta(L_{1},L_{2})$, we measure the depth $\alpha_{\text{depth}}(L_{1},L_{2})$ between $L_{1}$ and $L_{2}$ as:
$\alpha_{\mathrm{depth}}=1-\frac{\delta(L_{1},L_{2})}{\max(\text{depth}(L_{1}),\text{depth}(L_{2}))}$ (7)
The final language similarity score is computed as:
$\mathrm{Sim}(L_{1},L_{2})=S_{\mathrm{distance}}\times\alpha_{\mathrm{depth}}$ (8)
Appendix B Appendix: Neuron Patterns
B.1 Deactivation Overlap Neurons
Figure 7 presents overlap neurons distributions and their deactivation effects on Chinese-Hebrew BLI task.
Figure 7: Overlap neurons distributions and their deactivation effects on Chinese-Hebrew BLI task.
B.2 Linguistic Spectrum in Mistral
Figure 8: Mistral’s linguistic spectrum across 15 languages from 7 families. The color intensity represents the degree of overlap between language pairs.
B.3 Parameter $K$ Discussion
Here we discuss which $k$ middle layers should be selected to quantify linguistic similarity. According to observations in section 3.3.3, neurons in middle-layers should be prioritized over final-layers when measuring language similarity. We use embedding semantic similarity metric to determine $K$. We compute the layer-wise embedding semantic similarity between Arabic-Hebrew and Chinese-Hebrew pairs when predicting the same Hebrew token. As presented in Figure 9, we find embedding similarity is stable in the middle layers 10-21 and is not sensitive to variations in the predicted tokens. Therefore, $K$ layers is set to be 10-21 in LLaMA 3 and 15-23 in Mistral.
Figure 9: Embedding semantic similarity between Arabic-to-Hebrew and Chinese-to-Hebrew translations when predicting the same token at each layer.
B.4 Overlap Neuron Distribution
Figure 10 presents the distribution of overlap neurons across different language pairs in LLaMA 3.
Figure 10: Distribution of overlap neurons across different language pairs in LLaMA 3.
Appendix C Appendix: Ablation Results
This section presents the detailed experimental ablation to evaluate the impact of neuron probe data construction and the HSIC similarity metric on bridge selection. Table 3 and Table 4 present the results of ablation experiments on the BLI task. “w/o *” denotes replacing our probe data with a simplified version based on FLORES+. For example, “w/o *” probe data in Indonesian-Hebrew is illustrated in Figure 11.
Table 4: Performance comparison of using HSIC and Cosine similarity metrics on the BLI task.
LLaMA-3-8B
Bridge HSIC Bridge Cosine
Zh-Ja En 77.90 $\uparrow$ De 76.60
Zh-He En 56.70 $\downarrow$ De 58.60
Zh-Tl En 54.00 $\uparrow$ De 51.00
Zh-Sw En 36.00 $\uparrow$ De 33.80
Ar-Ja En 80.60 $\uparrow$ De 76.30
Ar-He En 64.50 $\uparrow$ De 64.10
Ar-Tl En 56.80 $\uparrow$ De 52.80
Ar-Sw En 42.40 $\downarrow$ De 43.00
Id-Ja En 66.10 $\uparrow$ De 65.10
Id-He Es 61.30 $\uparrow$ De 59.90
Id-Tl Fr 61.00 $\uparrow$ De 57.00
Id-Sw Fr 30.00 $\uparrow$ De 28.30
En-He Es 71.80 $\downarrow$ Pt 72.30
En-Tl Fr 58.70 $\uparrow$ Pt 55.80
En-Sw Fr 26.20 $\uparrow$ It 21.50
Figure 11: Example of “w/o *” probe data in Indonesian-Tagalog.
Appendix D Appendix: Neuron Semantic Analysis
In this section, we analyze the semantic similarity of overlapping neurons in two language groups: Hebrew-Tagalog-Swahili (He-Tl-Sw, different language family) and Portuguese-Spanish-Italian (pt-Es-It, same language families). For each group, we select overlapping neurons and the same number of randomly sampled neurons for comparison. We input $m$ parallel sentences and record the neuron activation frequency for each sentence, obtaining three $m\times 100$ activation matrices for overlap and random neurons. These matrices are mapped to a 2D semantic space using UMAP McInnes et al. (2018), with each point representing a neuron activated by a sentence.
As shown in Figure 12, randomly sampled neurons align with linguistic relationships. For example, random neurons of He and Tl are close, indicating semantic proximity in the same family. The overlap neurons cluster together, both within and across language families, proving that our approach can effectively capture semantic similarity.
Figure 12: Visualization of semantic similarity comparing language-overlapping neurons and randomly sampled neurons.
Appendix E Appendix: Prompt Templates